# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ A.2 Experiments in Hidden Preferences Environment
Experimental protocol. Given that PRUCB-HP models both the expected arms reward and user preference, we designed a new user-switching protocol for evaluation. Figure 7 illustrates this protocol with 3 users and 9 arms. Specifically, at each episode, one user is exposed to a block of arms (3 in our illustration). Only the arms within this block can be selected for this user. After one arm has been pulled, the system observes the reward $\boldsymbol{r}_{a_{t},t}$ and user’s overall ratings $g_{a_{t},t}$ corresponding to the pulled arm $a_{t}$. In the next episode, the arm block rotates to another user. The goal is to maximize the cumulative overall ratings from all users.
@@
Baselines. For performance comparison, we choose the MO-MAB baselines used in unknown environment (Appendix A.1, including S-UCB [1], S-MOSS, Pareto-UCB [1] and Pareto-TS [13]). Besides, we develop a variant of OFUL [19], a widely used linear bandit benchmark, for comparison. This variant estimates user preferences using ridge regression based on reward and overall reward feedback, replaces the input feature with the empirical reward estimate, and applies an $\epsilon=0.05$ rate for exploration. Additionally, note that the scale overall score is also provided, it is feasible to use standard MAB methods by leveraging historical overall rewards for optimization. Hence we also choose classic MAB algorithms including UCB [17] and MOSS [23] for comparison.
Experimental settings. In our experiment, we set $N$ users and $3N$ arms in total, and each arm associates with $D$-dimensional reward. The generations of instantaneous reward $\boldsymbol{r}_{i,t}$ of arms and user preference $\boldsymbol{c}_{t}$ follow the same settings as unknown environment.
@@
Implementation. Similarly, we set $\alpha=1$ in PRUCB-HP. For regularization coefficient, we set $\lambda=1$. For confidence radius, we set $\beta_{t}=0.1\sqrt{D\log(t)}$. For weight scalar, we set $\omega=D$. We perform 10 trials up to round $T=5000$ for each set of evaluation.
Results. We report average performance of the algorithms in Fig. 8. As shown, our proposed PRUCB-HP achieves superior results in terms of regret under all experimental settings compared to other competitors. This empirical evidence suggests that modeling user preference and leveraging this information for arm selection significantly enhances the performance of customized bandits optimization.
A.2.1 Analysis on WLS Estimator
In this section, we investigate the effectiveness of the proposed WLS-preference estimator with the carefully designed weight $w_{t}$. Specifically, we consider the 2D PAMO-MAB toy instance shown in Section 6.1. This instance contains two arms: Arm-1 with dominated mean reward $[0.2,0.2]^{\top}$, and Arm-2 with Pareto-optimal mean reward $[0.8,0.8]^{\top}$. The preference $\boldsymbol{c}_{t}$ at each step follows a Gaussian distribution with a mean of $[0.5,0.5]$ and variance of $0.05$ for each objective dimension.
@@ A.2.1 Analysis on WLS Estimator
Appendix B Proof of Proposition 1
Lemma 7 (Variant of Lemma 7 in [25]).
@@ Appendix B Proof of Proposition 1
The sub-linear regret bound implies that for a sufficiently large $T$ there exists a constant $C>0$ such that $\sum_{i=1}^{K}\mathbb{E}[N_{i,T}]\boldsymbol{\overline{c}}_{t}^{T}(\mu_{a^{*}}-\mu_{i})<CT$. Hence we have $\mathbb{E}[N_{i,T}]\boldsymbol{\overline{c}}_{t}^{T}(\mu_{a^{*}}-\mu_{i})\leq CT,\forall i\neq a^{*}$, implying $\mathbb{E}[N_{i,T}]<\frac{CT}{\boldsymbol{\overline{c}}_{t}^{T}(\mu_{a^{*}}-\mu_{i})}$. ∎
@@
We proof the Proposition 1 using a simple toy instance with 2 arms and 2 objectives. For the derivation based on the general instance, please refer to Appendix B.1.
@@
Assume there exists a preferences-free algorithm $\mathcal{A}$ (i.e., Pareto-UCB [1]) achieving sub-linear regret under preference $\boldsymbol{c}_{1}$. By Lemma 7, we have
@@
B.1 General Version
Definition 2 (Pareto order, [3]).
@@ B.1 General Version
Due to $\boldsymbol{\mu}_{u}\neq\boldsymbol{\mu}_{v}$, by taking the orthogonal complement of $\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}$, we can construct a subset $\mathcal{C}_{\varsigma^{+}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{T}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=0\}$. Next we consider two different constant preferences vector sets as the user’s preferences, to construct two sets of preferences-aware MO-MAB scenarios.
@@
Scenarios $\mathcal{S}_{\varsigma^{+}}$. For any $\varsigma^{+}>0$, we can construct a subset $\mathcal{C}_{\varsigma^{+}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{T}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=\varsigma^{+}\}$. Specifically, the general form of $\boldsymbol{c}_{\varsigma^{+}}\in\mathcal{C}_{\varsigma^{+}}$ can be written as $\boldsymbol{c}_{\varsigma^{+}}=\frac{\varsigma^{+}}{\|\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}\|_{2}^{2}}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})+\boldsymbol{c}_{0}$, where $\boldsymbol{c}_{0}$ is any vector such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$. Then for the preferences-aware MO-MAB scenarios $\mathcal{S}_{\varsigma^{+}}:=\{\mathcal{P}\times\mathcal{C}_{\varsigma^{+}}\}$ under the sets of arm reward distributions $\mathcal{P}$ and user preferences $\mathcal{C}_{\varsigma^{+}}$, it is obvious that arm $u$ is the optimal arm since $\boldsymbol{\mu}_{u}\succ\boldsymbol{\mu}_{i},\forall i\in[K]\setminus\{u,v\}$ and $\boldsymbol{c}_{\varsigma^{+}}^{T}\boldsymbol{\mu}_{u}>\boldsymbol{c}_{\varsigma^{+}}^{T}\boldsymbol{\mu}_{v},\forall\boldsymbol{c}_{\varsigma^{+}}\in\mathcal{C}_{\varsigma^{+}}$.
@@
Scenarios $\mathcal{S}_{\varepsilon^{-}}$. Similarly, for any $\varepsilon^{-}<0$, we can construct a subset $\mathcal{C}_{\varepsilon^{-}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{T}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=\varepsilon^{-}\}$, with the general form of $\boldsymbol{c}_{\varepsilon^{-}}=\frac{\varepsilon^{-}}{\|\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}\|_{2}^{2}}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})+\boldsymbol{c}_{0}$, where $\boldsymbol{c}_{0}$ is any vector such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$. For scenarios $\mathcal{S}_{\varepsilon^{-}}:=\{\mathcal{P}\times\mathcal{C}_{\varepsilon^{-}}\}$ with same arm rewards distributions $\mathcal{P}$ but modified user preferences $\mathcal{C}_{\varepsilon^{-}}$ sets, we have the arm $v$ to be the optimal.
@@
We use $\mathbb{P}_{\varsigma^{+}}$ to denote the probability with respect to the scenarios $\mathcal{S}_{\varsigma^{+}}$, and use $\mathbb{P}_{\varepsilon^{-}}$ to denote the probability conditioned on $\mathcal{S}_{\varepsilon^{-}}$. Analogous expectations $\mathbb{E}_{\varsigma^{+}}[\cdot]$ and $\mathbb{E}_{\varepsilon^{-}}[\cdot]$ will also be used. Let $\boldsymbol{\mathrm{a}}^{t-1}=\{A_{1},...,A_{t-1}\}$ and $\boldsymbol{\mathrm{r}}^{t-1}=\{\boldsymbol{x}_{1},...,\boldsymbol{x}_{t-1}\}$ be the actual sequence of arms pulled and the sequence of received rewards up to episode $t-1$, and $\boldsymbol{\mathrm{H}}^{t-1}=\{\langle A_{1},\boldsymbol{x}_{1}\rangle,...,\langle A_{t-1},\boldsymbol{x}_{t-1}\rangle\}$ be the corresponding historical rewards sequence. For consistency, we define $\boldsymbol{\mathrm{a}}^{0}$, $\boldsymbol{\mathrm{r}}^{0}$ and $\boldsymbol{\mathrm{H}}^{0}$ as the empty sets. Assume there exists a preferences-free algorithm $\mathcal{A}$ (i.e., Pareto-UCB [1]) that is possibly dependent on historical rewards sequence $\boldsymbol{\mathrm{H}}^{t-1}$ at episode $t$ (classical assumption in MAB), achieving sub-linear regret in scenarios $\mathcal{S}_{\varsigma^{+}}$. Let $N_{i,T}$ be the number of pulls of arm $i$ by $\mathcal{A}$ up to $T$ episode. By Lemma 7, we have
@@
Since the policy $\pi^{\mathcal{A}}_{t}$ of $\mathcal{A}$ is possibly dependent on $\boldsymbol{\mathrm{H}}^{t-1}$ but independent on the sequences of instantaneous preferences $\boldsymbol{\mathrm{c}}^{t}$ and preferences means $\boldsymbol{\overline{c}}$, for $t\in(0,T\}$, $i\in[K]$ we have
@@
[…]
@@
where $\boldsymbol{c}_{0},\boldsymbol{c}_{\varepsilon^{-}}$ can be any constant vectors such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$ and $\boldsymbol{c}_{0}\in\mathcal{C}_{\varepsilon^{-}}$. (a) holds since the policy $\pi^{\mathcal{A}}_{t}$ is independent of $\boldsymbol{\mathrm{c}}^{t}$ and $\boldsymbol{\overline{c}}$. Hence $\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{0}]^{t},\boldsymbol{c}_{0})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{\varepsilon^{-}}]^{t},\boldsymbol{c}_{\varepsilon^{-}})$ (recall the definition of preferences-free algorithm in Definition 1).
@@
[…]
@@
$\mathbb{E}_{\varsigma^{+}}[N_{i,T}]=\sum_{t=1}^{T}\mathbb{E}_{\varsigma^{+}}[\mathds{1}_{a_{t}=i}]=\sum_{t=1}^{T}\mathbb{E}_{\varepsilon^{-}}[\mathds{1}_{a_{t}=i}]=\mathbb{E}_{\varepsilon^{-}}[N_{i,T}].$
@@
$R(T)$ $=\sum_{i\neq v}\boldsymbol{c}_{\varepsilon^{-}}^{T}(\mu_{v}-\mu_{i})\mathbb{E}_{\varepsilon^{-}}[N_{i,T}]$

### v2

@@ B.1 Experiments in Hidden Preferences Environment
Experimental protocol.
Given that PRUCB-HP models both the expected arms reward and user preference, we designed a new user-switching protocol for evaluation. Figure 6 illustrates this protocol with 3 users and 9 arms. Specifically, at each episode, one user is exposed to a block of arms (3 in our illustration). Only the arms within this block can be selected for this user. After one arm has been pulled, the system observes the reward $\boldsymbol{r}_{a_{t},t}$ and user’s overall ratings $g_{a_{t},t}$ corresponding to the pulled arm $a_{t}$. In the next episode, the arm block rotates to another user. The goal is to maximize the cumulative overall ratings from all users.
@@
Baselines.
We compare our algorithm PRUCB-HP in terms of regret with the following multi-objective bandits algorithms.
  • S-UCB (Drugan and Nowe 2013): the scalarized UCB algorithm, which scalarizes the multi-dimensional reward by assigning weights to each objective and then employs the single objective UCB algorithm Auer, Cesa-Bianchi, and Fischer (2002). Throughout the experiments, we assign each objective with equal weight.
  • S-MOSS: the scalarized UCB algorithm, which follows the similar way with S-UCB by scalarizing the multi-dimensional reward into a single one, but uses MOSS (Audibert and Bubeck 2009) policy for arm selection.
  • Pareto-UCB (Drugan and Nowe 2013): the Pareto-based algorithm, which compares different arms by the upper confidence bounds of their expected multi-dimensional reward by Pareto order and pulls an arm uniformly from the approximate Pareto front.
  • Pareto-TS (Yahyaa and Manderick 2015): the Pareto-based algorithm, which makes use of the Thompson sampling technique to estimate the expected reward for every arm and selects an arm uniformly at random from the estimated Pareto front.
  • OFUL (Abbasi-Yadkori, Pál, and Szepesvári 2011): an adaptation variant of OFUL (Abbasi-Yadkori, Pál, and Szepesvári 2011), a widely used linear bandit benchmark, in MO-MAB problem. This variant estimates user preferences using ridge regression based on reward and overall reward feedback, replaces the input feature with the empirical reward estimate, and applies an $\epsilon=0.05$ rate for exploration.
  • UCB (Auer, Cesa-Bianchi, and Fischer 2002), MOSS (Audibert and Bubeck 2009): classic algorithms for single-objective Multi-Armed Bandits (MAB) that rely on feedback from overall scalar rewards.
Experimental settings.
For evaluation, we use a synthetic dataset. Specifically, we consider the MO-MAB with $N$ users, $K$ arms, each arm $i\in[K]$ associated with a $D$-dimensional reward, where the reward of each objective $d$ follows a Gaussian distribution with a randomized mean $\boldsymbol{\mu}_{i}(d)\in[0,1]$ and variance of 0.01. For each user $n\in[N]$ preference, we consider two settings including predefined preference and randomized preference. For predefined preference-aware structure, we define the mean preference $\boldsymbol{\overline{c}}^{n}$ as $\boldsymbol{\overline{c}}_{n}(d)=2.0\text{ if }d=j;0.5\text{ otherwise}$, where $j\in[D]$ is randomly selected. The practical implication of this structure is that it represents a common scenario in which the user exhibits a markedly higher preference for one particular objective while showing little interest in others. For randomized preference, the values of mean preference $\boldsymbol{\overline{c}}^{n}$ are randomly defined within $[0,5]$. For both setups, the instantaneous preference is generated under Gaussian distributions with corresponding means and variance of 0.5.
@@
Implementation.
Similarly, we set $\alpha=1$ in PRUCB-HP. For regularization coefficient, we set $\lambda=1$. For confidence radius, we set $\beta_{t}=0.1\sqrt{D\log(t)}$. For weight scalar, we set $\omega=D$. We perform 10 trials up to round $T=5000$ for each set of evaluation.
Regret Results.
We report average performance of the algorithms in Fig. 7. As shown, our proposed PRUCB-HP achieves superior results in terms of regret under all experimental settings compared to other competitors. This empirical evidence suggests that modeling user preference and leveraging this information for arm selection significantly enhances the performance of customized bandits optimization.
Analysis on WLS Estimator
In this section, we investigate the effectiveness of the proposed WLS-preference estimator with the carefully designed weight $w_{t}$. Specifically, we consider the 2D PAMO-MAB toy instance shown in Section 4.1. This instance contains two arms: Arm-1 with dominated mean reward $0.2,0.2]^{\top}$, and Arm-2 with Pareto-optimal mean reward $0.8,0.8]^{\top}$. The preference $\boldsymbol{c}_{t}$ at each step follows a Gaussian distribution with a mean of $[0.5,0.5]$ and variance of $0.05$ for each objective dimension.
@@
B.2 Experiments in Unknown Preference Environment
In this section, we verify the capability of PRUCB-UP to model users’ preferences $\boldsymbol{c}_{t}^{n}$ and optimize the overall reward cross all users in a stationary preference environment. We compare our algorithm PRUCB-UP in terms of regret with the following multi-objective bandits algorithms including S-UCB (Drugan and Nowe 2013), S-MOSS, Pareto-UCB (Drugan and Nowe 2013) and Pareto-TS (Yahyaa and Manderick 2015)). Additionally, we compare the performance with GT: a PRUCB-UP variant with known preference as a ground truth, where we replace the preference estimation in Algorithm 2 with the true preference estimation.
Experimental Settings and Implementations. The generations of user preference $\boldsymbol{c}_{t}^{n}$ and objective rewards $\boldsymbol{r}_{t}$ follow the same settings as hidden environment. For the implementations of the algorithms, following previous studies (Auer, Cesa-Bianchi, and Fischer 2002; Audibert, Munos, and Szepesvári 2007), we set $\alpha=1$. The time horizon is set to $T=5000$ rounds, the number of user $N$ is set as 2, and we repeat 10 trials for each set of evaluation due to the randomness from both environment and algorithms.
Regret Results. We report the averaged regret performance of the algorithms in Fig. 9. It is evident that our algorithm significantly outperform other competitors in all experiments. This is expected since the competing algorithms are designed for Pareto-optimality identification and do not utilize the preference structure of users considered in this paper, which our algorithm explicitly exploits. Additionally, from the zoom-in window, we observe that PRUCB-UP exhibits only a very slight performance degradation compared to GT, which knows the preference expectation in advance. This indicates that the proposed PRUCB-UP can effectively model user preference via empirical estimation in stationary preference environments.
Appendix C Proof of Proposition 1
Lemma 4 (Variant of Lemma 7 in (Jun et al. 2018)).
@@ Appendix C Proof of Proposition 1
The sub-linear regret bound implies that for a sufficiently large $T$ there exists a constant $C>0$ such that $\sum_{i=1}^{K}\mathbb{E}[N_{i,T}]\boldsymbol{\overline{c}}_{t}^{\top}(\mu_{a^{*}}-\mu_{i})<CT$. Hence we have $\mathbb{E}[N_{i,T}]\boldsymbol{\overline{c}}_{t}^{\top}(\mu_{a^{*}}-\mu_{i})\leq CT,\forall i\neq a^{*}$, implying $\mathbb{E}[N_{i,T}]<\frac{CT}{\boldsymbol{\overline{c}}_{t}^{\top}(\mu_{a^{*}}-\mu_{i})}$. ∎
@@
We proof the Proposition 1 using a simple toy instance with 2 arms and 2 objectives. For the derivation based on the general instance, please refer to Appendix C.1.
@@
Assume there exists a preferences-free algorithm $\mathcal{A}$ (i.e., Pareto-UCB (Drugan and Nowe 2013)) achieving sub-linear regret under preference $\boldsymbol{c}_{1}$. By Lemma 4, we have
@@
C.1 General Version
Definition 2 (Pareto order, (Lu et al. 2019)).
@@ C.1 General Version
Due to $\boldsymbol{\mu}_{u}\neq\boldsymbol{\mu}_{v}$, by taking the orthogonal complement of $\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}$, we can construct a subset $\mathcal{C}_{\varsigma^{+}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{\top}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=0\}$. Next we consider two different constant preferences vector sets as the user’s preferences, to construct two sets of preferences-aware MO-MAB scenarios.
@@
Scenarios $\mathcal{S}_{\varsigma^{+}}$. For any $\varsigma^{+}>0$, we can construct a subset $\mathcal{C}_{\varsigma^{+}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{\top}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=\varsigma^{+}\}$. Specifically, the general form of $\boldsymbol{c}_{\varsigma^{+}}\in\mathcal{C}_{\varsigma^{+}}$ can be written as $\boldsymbol{c}_{\varsigma^{+}}=\frac{\varsigma^{+}}{\|\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}\|_{2}^{2}}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})+\boldsymbol{c}_{0}$, where $\boldsymbol{c}_{0}$ is any vector such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$. Then for the preferences-aware MO-MAB scenarios $\mathcal{S}_{\varsigma^{+}}:=\{\mathcal{P}\times\mathcal{C}_{\varsigma^{+}}\}$ under the sets of arm reward distributions $\mathcal{P}$ and user preferences $\mathcal{C}_{\varsigma^{+}}$, it is obvious that arm $u$ is the optimal arm since $\boldsymbol{\mu}_{u}\succ\boldsymbol{\mu}_{i},\forall i\in[K]\setminus\{u,v\}$ and $\boldsymbol{c}_{\varsigma^{+}}^{\top}\boldsymbol{\mu}_{u}>\boldsymbol{c}_{\varsigma^{+}}^{\top}\boldsymbol{\mu}_{v},\forall\boldsymbol{c}_{\varsigma^{+}}\in\mathcal{C}_{\varsigma^{+}}$.
@@
Scenarios $\mathcal{S}_{\varepsilon^{-}}$. Similarly, for any $\varepsilon^{-}<0$, we can construct a subset $\mathcal{C}_{\varepsilon^{-}}:=\{\boldsymbol{c}\in\mathbb{R}^{D}|\boldsymbol{c}^{\top}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})=\varepsilon^{-}\}$, with the general form of $\boldsymbol{c}_{\varepsilon^{-}}=\frac{\varepsilon^{-}}{\|\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v}\|_{2}^{2}}(\boldsymbol{\mu}_{u}-\boldsymbol{\mu}_{v})+\boldsymbol{c}_{0}$, where $\boldsymbol{c}_{0}$ is any vector such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$. For scenarios $\mathcal{S}_{\varepsilon^{-}}:=\{\mathcal{P}\times\mathcal{C}_{\varepsilon^{-}}\}$ with same arm rewards distributions $\mathcal{P}$ but modified user preferences $\mathcal{C}_{\varepsilon^{-}}$ sets, we have the arm $v$ to be the optimal.
@@
We use $\mathbb{P}_{\varsigma^{+}}$ to denote the probability with respect to the scenarios $\mathcal{S}_{\varsigma^{+}}$, and use $\mathbb{P}_{\varepsilon^{-}}$ to denote the probability conditioned on $\mathcal{S}_{\varepsilon^{-}}$. Analogous expectations $\mathbb{E}_{\varsigma^{+}}[\cdot]$ and $\mathbb{E}_{\varepsilon^{-}}[\cdot]$ will also be used. Let $\boldsymbol{\mathrm{a}}^{t-1}=\{A_{1},...,A_{t-1}\}$ and $\boldsymbol{\mathrm{r}}^{t-1}=\{\boldsymbol{x}_{1},...,\boldsymbol{x}_{t-1}\}$ be the actual sequence of arms pulled and the sequence of received rewards up to episode t1t-1, and $\boldsymbol{\mathrm{H}}^{t-1}=\{\langle A_{1},\boldsymbol{x}_{1}\rangle,...,\langle A_{t-1},\boldsymbol{x}_{t-1}\rangle\}$ be the corresponding historical rewards sequence. For consistency, we define $\boldsymbol{\mathrm{a}}^{0}$, $\boldsymbol{\mathrm{r}}^{0}$ and $\boldsymbol{\mathrm{H}}^{0}$ as the empty sets. Assume there exists a preferences-free algorithm $\mathcal{A}$ (i.e., Pareto-UCB (Drugan and Nowe 2013)) that is possibly dependent on historical rewards sequence $\boldsymbol{\mathrm{H}}^{t-1}$ at episode $t$ (classical assumption in MAB), achieving sub-linear regret in scenarios $\mathcal{S}_{\varsigma^{+}}$. Let $N_{i,T}$ be the number of pulls of arm $i$ by $\mathcal{A}$ up to $T$ episode. By Lemma 4, we have
@@
Since the policy $\pi^{\mathcal{A}}_{t}$ of $\mathcal{A}$ is possibly dependent on $\boldsymbol{\mathrm{H}}^{t-1}$ but independent on the sequences of instantaneous preferences $\boldsymbol{\mathrm{c}}^{\top}$ and preferences means $\boldsymbol{\overline{c}}$, for $t\in(0,T\}$, $i\in[K]$ we have
@@
$\quad=\sum_{\begin{subarray}{c}\boldsymbol{\mathrm{a}}^{t-1}\in[K]^{t-1}\end{subarray}}\int_{\begin{subarray}{c}\boldsymbol{\mathrm{r}}^{t-1}\in[0,1]^{D\times(t-1)}\end{subarray}}\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{0}]^{\top},\boldsymbol{c}_{0})\cdot\mathbb{P}_{\varsigma^{+}}(\boldsymbol{\mathrm{H}}^{t-1})d\boldsymbol{\mathrm{r}}^{t-1}$
$\quad\quad\quad\quad-\sum_{\begin{subarray}{c}\boldsymbol{\mathrm{a}}^{t-1}\in[K]^{t-1}\end{subarray}}\int_{\begin{subarray}{c}\boldsymbol{\mathrm{r}}^{t-1}\in[0,1]^{D\times(t-1)}\end{subarray}}\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{\varepsilon^{-}}]^{\top},\boldsymbol{c}_{\varepsilon^{-}})\cdot\mathbb{P}_{\varepsilon^{-}}(\boldsymbol{\mathrm{H}}^{t-1})d\boldsymbol{\mathrm{r}}^{t-1}$
$\quad\underset{(a)}{=}\sum_{\begin{subarray}{c}\boldsymbol{\mathrm{a}}^{t-1}\in[K]^{t-1}\end{subarray}}\int_{\begin{subarray}{c}\boldsymbol{\mathrm{r}}^{t-1}\in[0,1]^{D\times(t-1)}\end{subarray}}\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1})\cdot\bigg(\mathbb{P}_{\varsigma^{+}}(\boldsymbol{\mathrm{H}}^{t-1})-\mathbb{P}_{\varepsilon^{-}}(\boldsymbol{\mathrm{H}}^{t-1})\bigg)d\boldsymbol{\mathrm{r}}^{t-1},$
@@
where $\boldsymbol{c}_{0},\boldsymbol{c}_{\varepsilon^{-}}$ can be any constant vectors such that $\boldsymbol{c}_{0}\in\mathcal{C}_{0}$ and $\boldsymbol{c}_{0}\in\mathcal{C}_{\varepsilon^{-}}$. (a) holds since the policy $\pi^{\mathcal{A}}_{t}$ is independent of $\boldsymbol{\mathrm{c}}^{\top}$ and $\boldsymbol{\overline{c}}$. Hence $\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{0}]^{\top},\boldsymbol{c}_{0})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1},[\boldsymbol{c}_{\varepsilon^{-}}]^{\top},\boldsymbol{c}_{\varepsilon^{-}})$ (recall the definition of preferences-free algorithm in Definition 1).
@@
$\quad=\sum_{\begin{subarray}{c}\boldsymbol{\mathrm{a}}^{t-1}\in[K]^{t-1}\end{subarray}}\int_{\begin{subarray}{c}\boldsymbol{\mathrm{r}}^{t-1}\in[0,1]^{D\times(t-1)}\end{subarray}}\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i|\boldsymbol{\mathrm{H}}^{t-1})\cdot\bigg(\cancelto{0}{\mathbb{P}_{\varsigma^{+}}(\boldsymbol{\mathrm{H}}^{t-1})-\mathbb{P}_{\varepsilon^{-}}(\boldsymbol{\mathrm{H}}^{t-1})}\bigg)d\boldsymbol{\mathrm{r}}^{t-1}=0.$
@@
$\mathbb{E}_{\varsigma^{+}}[N_{i,T}]=\sum_{t=1}^{\top}\mathbb{E}_{\varsigma^{+}}[\mathds{1}_{a_{t}=i}]=\sum_{t=1}^{\top}\mathbb{E}_{\varepsilon^{-}}[\mathds{1}_{a_{t}=i}]=\mathbb{E}_{\varepsilon^{-}}[N_{i,T}].$
@@
$R(T)$ $=\sum_{i\neq v}\boldsymbol{c}_{\varepsilon^{-}}^{\top}(\mu_{v}-\mu_{i})\mathbb{E}_{\varepsilon^{-}}[N_{i,T}]$